OpenStudy (anonymous):

If 8 sin t − cos t = 0, find all possible values of sin t and cos t

OpenStudy (anonymous):

do you mean find all possible values of t?

OpenStudy (dumbcow):

sin t = 1/8*cos t tan t = 1/8 t = 7.125 and 187.125

OpenStudy (cwrw238):

let 8 = R cos a and 1 = R sin a R sin (t - a) = 0 solve this sytem of equations

OpenStudy (mertsj):

dumbcow is correct except if you need ALL possible values of t, it would be 7.125 and 187.125 +360n